# DyWA: Dynamics-adaptive World Action Model for Generalizable Non-prehensile Manipulation

### v1

@@ 2.2 World Models
The world model [14] learns a generalized representation of the environment and predicts future world states based on sequences of actions. The development of world models has been extensively explored in domains such as gaming [13, 15, 16, 37] and autonomous driving [40, 19, 4]. Among these, the renowned Dreamer series [15, 16, 17] learns latent dynamics models from past experiences to predict state values and leverages these models for efficient actor learning.
In the domain of robotic manipulation, learning world models has also demonstrated significant effectiveness [10, 12, 3, 42]. A subset of these works [10, 12] trains dynamics models using offline robotic data and subsequently employs Model Predictive Control (MPC) during deployment. UniSim [42] directly learns a video generation model, which serves as a simulation environment for training manipulation policies. GR-2 [3] demonstrates that pre-training world models is critical for enhancing the generalization capability of downstream Vision-Language-Action (VLA) models during post-training.
Similar to JOWA [5], our world-action model simultaneously learns both the desired actions and the corresponding next states within a unified framework. Moreover, we incorporate dynamic conditions into the model, enabling the policy to adapt to and generalize across varying physical scenarios.
@@ 3.1 Task Formulation
Following HACMan and CORN, we focus on the task of 6D object rearrangement via non-prehensile manipulation. The robot’s objective is to execute a sequence of non-prehensile actions (i.e., pushing, flipping) to move an object on the table to a target 6D pose. We define the goal pose $G$ as a 6DoF transformation relative to the object’s initial pose, assuming both are stable on the table. The task state $S_{t}$ at timestep $t$ is represented by the relative transformation between the object’s current pose and the goal pose. The observation space includes a point cloud $P_{t}$ captured by a depth sensor, robot’s joint positions and velocities $J_{t}$, and the end-effector pose $E_{t}$ computed via forward kinematics.
@@ 3.2 Pipeline Overview
Our training pipeline follows a standard teacher-student distillation framework. Due to the difficulty of obtaining high-quality demonstrations for our task, we first train a state-based RL policy with additional privileged information—i.e., the full object point cloud, task state, and physical parameters—as the teacher policy. For consistency, we adopt the same reward design as CORN, as elaborated in the supplementary material. To obtain a vision-based policy suitable for real-world deployment, we introduce our Dynamics-conditioned World Action Model, which serves as the student policy and is designed to fully leverage the knowledge distilled from the teacher policy. Unlike the teacher, our student model relies solely on limited observations that are feasible to obtain in real-world settings.
In the following sections, we detail the design of the world action model (Sec. 3.3) and the dynamics adaptation mechanism (Sec. 3.4). To enable adaptive force interaction in this contact-rich task, we further incorporate a variable impedance controller as the action space (Sec. 3.5). Once trained (Sec. 3.6), our model can be transferred from simulation to the real world in a zero-shot manner, without requiring real-world fine-tuning.
@@ 3.3 World Action Model
Our model takes Observation and Goal Description as input, encoding different modalities using individual encoders. For the partial point cloud observation, we process it using a simplified PointNet++ [35] to obtain $f_{t}^{P}$, striking a balance between efficiency and capacity. The architectural details are provided in the supplementary material. For robot proprioception, we separately encode joint positions and velocities ($f_{t}^{J}$) and the end-effector pose ($f_{t}^{E}$) using shallow MLPs. For the Goal Description, instead of relying on the unknown task state $S_{t}$, we construct a visual goal representation by transforming the initial point cloud $P_{0}$ to the goal pose, yielding $P_{G}=GP_{0}$. This goal point cloud is then encoded using the same network as the observation point cloud encoder, ensuring consistency in feature extraction.
@@
We enforce the end-to-end model that jointly makes action decisions and predicts their outcomes, creating a synergistic learning process that, in turn, improves action learning. Specifically, the observation and goal embeddings are processed through MLPs to produce both the action $A_{t}$ and the next task state $S_{t+1}$, with supervision signals separately derived from the teacher policy and simulation outcomes. Our object-centric world model represents the environment using task state $S_{t+1}$ instead of high-dimensional visual signals, enabling the policy to focus on task-relevant dynamics. To represent rotations, we adopt the 9D representation [24, 27], and define the world model loss as:
$\mathcal{L}_{world}=\|T_{t+1}-\hat{T}_{t+1}\|^{2}+|\mathbf{R}_{t+1}-\hat{\mathbf{R}}_{t+1}|$ (1)
where $T_{t+1}\in\mathbb{R}^{3}$ and $\mathbf{R}_{t+1}\in SO(3)$ are the predicted translation and rotation, while $\hat{T}_{t+1}\in\mathbb{R}^{3}$ and $\hat{\mathbf{R}}_{t+1}\in SO(3)$ denote the ground-truth transformation obtained from simulation outcomes after action execution. Additionally, we employ an imitation loss, defined as the L2 loss between the predicted action and the teacher action:
$\mathcal{L}_{imitation}=\|A_{t}^{s}-A_{t}^{t}\|^{2}$ (2)
@@ 3.4 Dynamics Adaptation
$z_{t}=Embed(\{\text{concat}(f^{O}_{t-i-1},f^{A}_{t-i-2})\}_{i=1}^{L})$ (3)
@@
$\mathcal{L}_{adapt}=\|z_{t}^{Geo,Phy}-concat(f_{t}^{Geo},f_{t}^{Phy})\|^{2}$ (4)
@@ 3.5 Action Space with Variable Impedance
To enable adaptive force interaction between the robot and object, we employ variable impedance control as the low-level action execution mechanism. This allows the robot to dynamically regulate the interaction force based on task demands. Specifically, the action space of our policy consists of the subgoal residual of the end effector, $\Delta T_{ee}\in SE(3)$, along with joint-space impedance parameters. The joint-space impedance is parameterized by positional gains ($P\in\mathbb{R}^{7}$) and damping factors ($\rho\in\mathbb{R}^{7}$), where the velocity gains are computed as $D=\rho\sqrt{P}$. To execute the commanded end-effector motion, we first solve for the desired joint position using inverse kinematics with the damped least squares method [2]:
@@ 3.6 Training Protocol
$\mathcal{L}=\mathcal{L}_{imitation}+\mathcal{L}_{world}+\mathcal{L}_{adapt}$ (7)
We begin by training the teacher policy for 200K iterations in simulation using PPO. Subsequently, we employ DAgger to train the student policy under teacher supervision for 500K iterations. To enhance robustness and generalization, we introduce domain randomization during training by varying the object’s mass, scale, and friction, as well as the restitution properties of the object, table, and robot gripper. The object scale is adjusted such that its largest diameter remains within a predefined range. To further improve sim-to-real transfer, we inject small perturbations into the torque commands, object point cloud, and goal pose when training the student policy.
@@ 4 Experiments
Table 3: Quantitative results in the real world.
@@ 4.1 Benchmarking Tabletop Non-prehensile Rearrangement in Simulation
We evaluate our approach against two state-of-the-art baselines: HACMan and CORN, which represent primitive-based and closed-loop methods, respectively. Since HACMan was originally implemented in the MuJoCo simulator, we re-implemented it within our benchmark for a fair comparison. However, because it requires strict per-point correspondence as input, its success rate is extremely low in the unknown state setting. Therefore, we report its performance only on the easiest track. CORN shares the same simulation environment as our method, allowing us to train and evaluate it directly with minimal modifications. To ensure a fair comparison, we further enhanced CORN by replacing its shallow MLP-based point cloud encoder with the same vision backbone as ours. Additionally, for settings where the current object pose is unknown, we provided all methods with the same goal point cloud representation to maintain consistency.
@@ 4.2 Ablation Study
To analyze the optimization process, we visualize the loss curve during training and compare the approach that uses only dynamics adaptation (i.e., RMA) with that adding World Modeling. Our results show that during the distillation, simultaneous learning of the next state improves action coverage, confirming the synergy between world modeling and action learning. Additionally, we discuss the integration of the world model in the RL teacher policy, which is elaborated in the supplementary material.
Indivisibility of Dynamics Adaptation and World Modeling.
We investigate the individual and combined effects of dynamics adaptation and world modeling. Our results (Table 2) show that using only the world model or dynamics adaptation, i.e. RMA, provides only marginal improvements over the naive DAgger baseline, with success rates increasing by just 1.7% and 5.7%, respectively. However, when both modules are used together, the performance jumps significantly from 59.9% to 73.3%. This improvement can be attributed to the complementary nature of these components. Without dynamics adaptation, the world model lacks sufficient information to reason about the dynamic effects of interaction. Conversely, using only dynamics adaptation also provides limited benefits due to the absence of a sufficiently structured learning target. These findings highlight the indivisibility of world modeling and dynamics adaptation, demonstrating that their combination is a non-trivial yet highly effective design choice.
@@ 4.3 Real-World Experiments
To assess the effectiveness of dynamics adaptation, we conduct experiments on surfaces with varying friction coefficients. We select four tablecloths (Figure LABEL:fig:teaser) with progressive friction levels, i.e. $\mu_{1},\mu_{2},\mu_{3},\mu_{4}$ and use the bulldozer toy as the test object. Additionally, we report the average execution time for successful episodes. As shown in Table 4, the model without dynamics adaptation exhibits significant performance degradation when interacting with surfaces of different friction levels, leading to erratic execution times. In contrast, our policy with dynamics adaptation maintains consistent success rates while ensuring stable execution times across all surface conditions. This highlights the robustness of our approach in handling diverse real-world contact dynamics.
@@ 4.4 Applications
We present a practical manipulation system that integrates Vision-Language Models (VLMs), our non-prehensile policy, and a grasping model. By leveraging VLMs, our goal-conditioned policy can be executed based on natural language instructions. Specifically, we utilize SoFar [36], a model capable of generating semantic object poses from language commands, to specify goals for our policy. As shown in Figure 4, given the command “Put the grip of the electric drill into a person’s hand”, SoFar generates the target transformation of the drill (e.g., rotation $\Delta\theta=122^{\circ}$ and translation $\Delta x,\Delta y=[0.54,0.09]$), which is then used as the goal for our policy. This enables natural, instruction-driven object handovers, highlighting the potential of our approach in human-robot interaction.
Additionally, we showcase the system outperforms or complements traditional prehensile manipulation. As illustrated in the third row of Figure 3, a standard pick-and-place strategy struggles to flip a tiny switch due to gripper-table collisions, whereas our policy enables efficient rearrangement in a single continuous motion. Furthermore, when combined with an existing grasping model [38], our policy serves as an effective pre-grasping step. As shown in Figure 5, certain objects are inherently difficult to grasp due to their geometry—for example, a thin card lying flat on a surface or a broad cracker box exceeding the gripper’s maximum span. By leveraging a VLM-specified goal pose and applying our non-prehensile manipulation as a pre-grasping step, we can reorient these objects into grasp-friendly configurations, significantly improving grasp success rate.
@@ 5 Conclusion, Limitations, and Future Works
In this work, we present a novel policy learning approach that jointly predicts future states while adapting dynamics from historical trajectories. Our model enhances generalizable non-prehensile manipulation by reducing reliance on multi-camera setups and pose tracking modules while maintaining robustness across diverse physical conditions. Extensive simulation and real-world experiments validate the effectiveness of our approach. However, our method also has certain limitations since it relies solely on point clouds as the visual input modality. It struggles with symmetric objects due to geometric ambiguity, and faces challenges with transparent and specular objects, where raw depth is imcomplete. A promising direction is to incorporate additional appearance information to provide richer visual cues.

### v2

@@ 2.2 World Models
World models [13] learn compact representations of the environment and predict future states conditioned on action sequences. They have been widely studied in domains such as gaming [12, 14, 15, 44] and autonomous driving [47, 17, 4].
In robotic manipulation, jointly modeling future observations and actions has shown strong empirical performance [3, 50, 57, 48, 11]. We adopt the term world action model to characterize this class of policy learning methods with world model attribute, distinguishing it from prior work such as JOWA [5].
@@ 3.1 Task Formulation
Following HACMan and CORN, we focus on the task of 6D object rearrangement via non-prehensile manipulation. The robot’s objective is to execute a sequence of non-prehensile actions (i.e., pushing, flipping) to move an object on the table to a target 6D pose. We define the goal pose $\mathbf{G}$ as a 6DoF transformation relative to the object’s initial pose, assuming both are stable on the table. The task state $S_{t}$ at timestep $t$ is represented by the relative transformation between the object’s current pose and the goal pose. Observations include the partial point cloud $P_{t}$, joint states $J_{t}$, and end-effector pose $E_{t}$.
@@ 3.2 Pipeline Overview
Our training pipeline follows a standard teacher-student distillation framework. Due to the difficulty of obtaining high-quality demonstrations for our task, we first train a state-based RL policy with additional privileged information—i.e., the full object point cloud, task state, and physical parameters—as the teacher policy. For consistency, we adopt the same reward design as CORN, as elaborated in the supplementary material. To obtain a vision-based policy suitable for real-world deployment, we introduce our Dynamics-adaptive World Action Model, which serves as the student policy distilled from the teacher policy. Unlike the teacher, our student model relies solely on limited observations that are feasible to obtain in real-world settings.
In the following sections, we detail the design of the world action model (Sec. 3.3) and the dynamics adaptation mechanism (Sec. 3.4). To enable adaptive force interaction in this contact-rich manipulation, we further incorporate a variable impedance controller (Sec. 3.5). Once trained (Sec. 3.6), our model can be transferred from simulation to the real world in a zero-shot manner, without requiring real-world fine-tuning.
@@ 3.3 World Action Model
Definition.
A world action model refers to a policy model that jointly predicts actions and forecasts the corresponding future states. Although the current action is not provided as an explicit input, the model exhibits world model characteristics by implicitly conditioning on the current policy action prediction.
@@
Our model takes observation and goal description as input, encoding different modalities using individual encoders. For the partial point cloud observation, we process it using a simplified PointNet++ [41] to obtain $\mathbf{f}_{t}^{P}$. The architectural details are provided in the supplementary material. For robot proprioception, we separately encode joint positions and velocities ($\mathbf{f}_{t}^{J}$) and the end-effector pose ($\mathbf{f}_{t}^{E}$) using shallow MLPs. For the Goal Description, instead of relying on the unknown task state $S_{t}$, we construct a visual goal representation by transforming the initial point cloud $P_{0}$ to the goal pose, yielding $P_{G}=\mathbf{G}P_{0}$. This goal point cloud is then encoded using the shared network with the observation point cloud encoder.
@@
We enforce the end-to-end model that jointly makes action decisions and predicts their outcomes, creating a synergistic learning process that, in turn, improves action learning. Specifically, the observation and goal embeddings are processed through MLPs to produce both the action $\mathbf{A}_{t}$ and the next task state $S_{t+1}$, with supervision signals separately derived from the teacher policy and simulation outcomes. Our object-centric world model represents the environment using task state $S_{t+1}$ instead of high-dimensional visual signals, enabling the policy to focus on task-relevant dynamics. To represent rotations, we adopt the 9D representation [22, 25], and define the world model loss as:
$\mathcal{L}_{\text{world}}=\|\mathbf{T}_{t+1}-\hat{\mathbf{T}}_{t+1}\|_{2}^{2}+\|\mathbf{R}_{t+1}-\hat{\mathbf{R}}_{t+1}\|_{1}$ (1)
where $\mathbf{T}_{t+1}\in\mathbb{R}^{3}$ and $\mathbf{R}_{t+1}\in SO(3)$ are the predicted translation and rotation, while $\hat{\mathbf{T}}_{t+1}\in\mathbb{R}^{3}$ and $\hat{\mathbf{R}}_{t+1}\in SO(3)$ denote the ground-truth transformation obtained from simulation outcomes after action execution. Additionally, we employ an imitation loss, defined as the L2 loss between the predicted action and the teacher action:
$\mathcal{L}_{\text{imitation}}=\|\mathbf{A}_{t}^{s}-\mathbf{A}_{t}^{t}\|^{2}$ (2)
@@ 3.4 Dynamics Adaptation
$\mathbf{z}_{t}=\mathrm{Embed}\left(\left[\mathrm{concat}(\mathbf{f}^{O}_{t-i-1},\mathbf{f}^{A}_{t-i-2})\right]_{i=1}^{L}\right)$ (3)
@@
$\mathcal{L}_{\text{adapt}}=\|\mathbf{z}_{t}^{Geo,Phy}-concat(\mathbf{f}_{t}^{Geo},\mathbf{f}_{t}^{Phy})\|^{2}$ (4)
@@ 3.5 Action Space with Variable Impedance
To enable adaptive force interaction between the robot and object, we employ variable impedance control [8] as the low-level action execution mechanism. This allows the robot to dynamically regulate the interaction force based on task demands. Specifically, the action space of our policy consists of the subgoal residual of the end effector, $\Delta T_{ee}\in SE(3)$, along with joint-space impedance parameters. The joint-space impedance is parameterized by positional gains ($P\in\mathbb{R}^{7}$) and damping factors ($\rho\in\mathbb{R}^{7}$), where the velocity gains are computed as $D=\rho\sqrt{P}$. To execute the commanded end-effector motion, we first solve for the desired joint position using inverse kinematics with the damped least squares method [2]:
@@ 3.6 Training Protocol
$\mathcal{L}=\mathcal{L}_{\text{imitation}}+\mathcal{L}_{\text{world}}+\mathcal{L}_{\text{adapt}}$ (7)
We begin by training the teacher policy for 200K iterations in simulation using PPO. Subsequently, we employ DAgger [43] to train the student policy under teacher supervision for 500K iterations. To enhance robustness and generalization, we introduce domain randomization during training by varying the object’s mass, scale, and friction, as well as the restitution properties of the object, table, and robot gripper. The object scale is adjusted such that its largest diameter remains within a predefined range. To further improve sim-to-real transfer, we inject small perturbations into the torque commands, object point cloud, and goal pose when training the student policy.
@@ 4 Experiments
Table 3: Quantitative results in the real world. Each cell shows the number of successful trials out of 5 attempts. Our method consistently achieves high success rates across diverse objects.
@@ 4.1 Benchmarking Tabletop Non-prehensile Rearrangement in Simulation
We evaluate our approach against two state-of-the-art baselines: HACMan and CORN, which represent primitive-based and closed-loop methods, respectively. Since HACMan was originally implemented in the MuJoCo simulator, we re-implemented it within our benchmark for a fair comparison. However, because it requires strict per-point correspondence as input, its success rate is extremely low in the unknown state setting. CORN shares the same simulation environment as our method, allowing us to train and evaluate it directly with minimal modifications. To ensure a fair comparison, we further enhanced CORN by replacing its shallow MLP-based point cloud encoder with the same vision backbone as ours. Additionally, for settings where the current object pose is unknown, we provided all methods with the same goal point cloud representation to maintain consistency.
@@ 4.2 Ablation Study
To analyze the optimization process, we visualize the loss curve during training and compare the approach that uses only dynamics adaptation (i.e., RMA) with that adding World Modeling. Our results show that during the distillation, simultaneous learning of the next state improves action loss convergence, confirming the synergy between world modeling and action learning. Additionally, we discuss the integration of the world model in the RL teacher policy, which is elaborated in the supplementary material.
On the Complementarity of Dynamics Adaptation and World Modeling.
We investigate the individual and combined effects of dynamics adaptation and world modeling. Our results (Table 2) show that using only the world model or dynamics adaptation, i.e. RMA, provides only marginal improvements over the naive DAgger baseline, with success rates increasing by just 1.7% and 5.7%, respectively. However, when both modules are used together, the performance jumps significantly from 59.9% to 73.3%. This improvement can be attributed to the complementary nature of these components. Without dynamics adaptation, the world model lacks sufficient information to reason about the dynamic effects of interaction. Conversely, using only dynamics adaptation also provides limited benefits due to the absence of a sufficiently structured learning target. These findings highlight the complementarity of world modeling and dynamics adaptation, demonstrating that their combination is a non-trivial yet highly effective design choice.
@@ 4.3 Real-World Experiments
To assess the effectiveness of dynamics adaptation, we conduct experiments on surfaces with varying friction coefficients. We select four tablecloths (Figure 6) with progressive friction levels, i.e. $\mu_{1},\mu_{2},\mu_{3},\mu_{4}$ and use the bulldozer toy as the test object. Additionally, we report the average execution time for successful episodes. As shown in Table 4, the model without dynamics adaptation exhibits significant performance degradation when interacting with surfaces of different friction levels, leading to erratic execution times. In contrast, our policy with dynamics adaptation maintains consistent success rates while ensuring stable execution times across all surface conditions. This highlights the robustness of our approach in handling diverse real-world contact dynamics.
@@ 4.4 Applications
We present a practical manipulation system that integrates Vision-Language Models (VLMs), our non-prehensile policy, and a grasping model [45]. By leveraging VLMs, our goal-conditioned policy can be executed based on natural language instructions. Specifically, we utilize SoFar [42], a model capable of generating semantic object poses from language commands, to specify goals for our policy. As shown in Figure 4, given the command “Put the grip of the electric drill into a person’s hand”, SoFar generates the target transformation of the drill (e.g., rotation $\Delta\theta=122^{\circ}$ and translation $\Delta x,\Delta y=[0.54,0.09]$), which is then used as the goal for our policy. This enables natural, instruction-driven object handovers, highlighting the potential of our approach in human-robot interaction.
Additionally, we showcase the system outperforms or complements traditional prehensile manipulation. As illustrated in the third row of Figure 3, a standard pick-and-place strategy struggles to flip a tiny switch due to gripper-table collisions, whereas our policy enables efficient rearrangement in a single continuous motion. Furthermore, our policy serves as an effective pre-grasping step in the system. As shown in Figure 5, certain objects are inherently difficult to grasp due to their geometry—for example, a thin card lying flat on a surface or a broad cracker box exceeding the gripper’s maximum span. Our system can firstly reorient these objects into grasp-friendly configurations, significantly improving grasp success rate.
@@ 5 Conclusion, Limitations, and Future Works
In this work, we present a novel policy learning approach that jointly predicts future states while adapting dynamics from historical trajectories. Our model enhances generalizable non-prehensile manipulation by reducing reliance on multi-camera setups and pose tracking modules while maintaining robustness across diverse physical conditions. Extensive simulation and real-world experiments validate the effectiveness of our approach. However, our method also has certain limitations since it relies solely on point clouds as the visual input modality. It struggles with symmetric objects due to geometric ambiguity, and faces challenges with transparent and specular objects, where raw depth is incomplete. A promising direction is to incorporate additional appearance information [29, 31, 30, 32, 27, 28, 26] to provide richer visual cues.
Acknowledgements
We thank Yixin Zheng for organizing the code release, and Junhao Yang for assistance with rendering. We also appreciate the valuable suggestions and discussions from Jiayi Chen, Jiazhao Zhang, Mi Yan and Shenyuan Gao. This work was supported in part by National Science and Technology Major Project (2022ZD0114904) & NSFC-6247070125.
Supplementary Material
Figure 6: Illustration of the high-level idea and generalization ability of DyWA. Given a target object’s 6D pose and single-view object point cloud, our non-prehensile manipulation policy aims to rearrange the object without grasping. Left: Our key insight is to enhance action learning by jointly predicting future states while adapting to dynamics from historical trajectories. (For clarity, rendered images are used for visualization, while the actual visual input consists of partial point clouds.) Right: After being trained in simulation, our policy achieves zero-shot sim-to-real transfer and generalizes across diverse dynamic properties, including variations in object geometry, object physical property (e.g., slipperiness and non-uniform mass distribution), and surface friction.
6 Real-world Setup
Figure 7: Real-world Experiment Setup
Figure 8: Objects used in the real-world experiments
As shown in Figure 7, We use a franka panda robot arm for execution and a Intel RealSense D435 camera for capturing point cloud. We also adopt the same object segmentation method with CORN [8], consists of color-based segmentation followed by depth-based back-projection to obtain a single-view object point cloud.
7 RL-based Teacher Policy
7.1 Reward Design
Following Kim et al. [19], Cho et al. [8], the reward function in our domain is defined as:
$r=r_{suc}+r_{reach}+r_{contact}-c_{energy},$ (8)
where $r_{suc}$ is the task success reward, $r_{reach}$ is the goal-reaching reward, $r_{contact}$ is the contact-inducing reward, and $c_{energy}$ is the energy penalty.
The task success reward is defined as:
$r_{suc}=\mathbbm{1}_{suc},$ (9)
where $\mathbbm{1}_{suc}$ is an indicator function that returns 1 when the object’s pose is within 0.05m and 0.1 radians of the target pose.
To facilitate learning, we introduce dense rewards $r_{reach}$ and $r_{contact}$, formulated based on a potential function [37] as:
$r=\gamma\phi(s^{\prime})-\phi(s),$ (10)
where $\gamma\in[0,1)$ is the discount factor. Specifically,
$\phi_{reach}(s)=k_{g}\gamma^{k_{d}\cdot d_{o,g}(s)},$ (11)
$\phi_{contact}(s)=k_{r}\gamma^{k_{d}\cdot d_{h,o}(s)},$ (12)
where $k_{g},k_{d},k_{r}\in\mathbb{R}$ are scaling coefficients. The term $d_{o,g}(s)$ represents the distance between the current object pose and the goal pose, measured using a bounding-box-based distance metric, while $d_{h,o}(s)$ denotes the distance between the object’s center of mass and the tip of the gripper.
The energy penalty is defined as:
$c_{energy}=k_{e}\sum_{i=1}^{7}\tau_{i}\dot{q}_{i},$ (13)
where $k_{e}\in\mathbb{R}$ is a scaling coefficient, and $\tau_{i}$ and $\dot{q}_{i}$ denote the torque and velocity of the $i^{\text{th}}$ joint, respectively.
7.2 Architecture
World Model FiLM Success Rate
94.1
93.9
93.5
Table 5: Success rate of RL-based Teacher Policy.
Our teacher policy architecture consists of separate encoders for each modality and an MLP-based policy network, which proves sufficiently effective for RL training. We also experiment with incorporating our proposed world modeling and FiLM into the teacher policy, as shown in Table 5. However, the performance remains nearly unchanged compared to the baseline approach. Together with the analysis in Figure 2, we attribute this to the RL policy already reaching its performance upper bound given the current architecture. The core contribution of DyWA primarily benefits the distillation process, facilitating better optimization of the imitation loss rather than improving the teacher policy itself.
8 Alternative of Dynamics Factor Conditioning
MLP Cross Atten FiLM
Success Rate 73.3 70.1 82.2
Table 6: Success Rate of student policy with different dynamics conditioning methods.
To investigate alternative conditioning mechanisms, we experimented with cross-attention layers as a replacement for FiLM. However, this approach led to significantly degraded performance. We hypothesize that the transformer-based cross-attention mechanism is more sensitive to data distribution shifts and may require additional architectural modifications or extensive data augmentation, introducing unnecessary overhead for this task. These findings further support FiLM as a lightweight yet effective method for integrating adaptation embeddings into the world-action model.
9 Simulation Assets and Setup
Figure 9: Simulation Environment Setup.
Figure 10: Objects used for training in the simulation benchmark.
Figure 11: Unseen objects used for evaluation in the simulation benchmark.
[…]
Hyperparameter Value
Learning rate 6e-4
Num. Environment 1024
Optimizer Adam
Normalization Layernorm
Dropout 0
Table 7: Hyper-parameters for Student’s Training Algorithm
Hyperparameter Value
Input Size (512, 3)
Key points $\mathbf{C_{i}}$ [64, 16]
Grouping Neighbours $\mathbf{K}$ 32
Grouped feature $\mathbf{M_{i}}$ [32, 128]
Global points MLP MLP(4096, 1024, 1024, 4096)
History length 5
History Decoder Conv1d+MaxPool
History Decoder channel 128
FiLM block Num 3
Pose Predictor Shared MLP(4096, 256)
Translation predictor MLP(256, 128, 64, 3)
Rotation predictor MLP(256, 128, 64, 3)
Actor MLP(4736, 1024, 256)
Table 8: Hyper-parameters for Student’s Encoder and Policy
Hyperparameter Value
RL algorithm PPO
Adam stepsize 3e-4
Num. Environment 4096
GAE parameter 0.95
Discount Factor 0.99
PPO clip range 0.3
Num. epoch 8
Table 9: Hyper-parameters for Teacher’s PPO Algorithm
Hyperparameter Value
Key points $\mathbf{C_{i}}$ [16]
Grouping Neighbors $\mathbf{K}$ 32
Grouped feature Channels $\mathbf{M_{i}}$ [128]
Shared MLP MLP(512, 256, 128)
Actor MLP(64, 20)
Critic MLP(64, 1)
Table 10: Hyper-parameters for Teacher’s Encoder and Policy
Figure 12: Qualatative results in the simulation.
10 Vision Encoder
We use a simplified version of PointNet++ [41] as our point cloud encoder. Specifically, the student policy encoder employs two layers of fixed-scale grouping, while the teacher policy encoder uses one layer. In the $\mathbf{i^{th}}$ grouping layer, $\mathbf{C_{i}}$ key points are selected via farthest point sampling (FPS), and each key point forms a group with its $\mathbf{K}$ nearest neighbors (KNN). Each cluster is then processed by two per-point MLP layers and two global MLP layers to generate a group feature. The output point cloud consists only of the $\mathbf{C_{i}}$ selected key points, each enriched with its corresponding $\mathbf{M_{i}}$-dimensional group feature. After the grouping layers, the per-point features are concatenated and passed through residual MLP layers, following the structure of PointNet++. The final output consists of several point tokens with grouped features.
11 Hyper-parameters
The following Tables 7, 8, 9, 10 demonstrate the hyper-parameters of our policy and the teacher policy.